OpenStudy (anonymous):

the news is on for 2 1/2 hours in the morning. 1/2 an hour at noon. 1/2 an hour in the evening. 1/2 an hour at night. how long is the news each day.

OpenStudy (anonymous):

help

OpenStudy (anonymous):

Add up all. \[2\frac{1}{2} + \frac{1}{2} +\frac{1}{2} +\frac{1}{2} = \frac{5}{2}+\frac{1}{2}+\frac{1}{2}+\frac{1}{2} = \frac{8}{2} = 4 \ hours\]